OpenStudy (anonymous):

The length of the top of the table is 7 m greater than the width. The area is 78m^2. Find the dimensions to the table.

OpenStudy (anonymous):

length = 13m width = 6m

OpenStudy (anonymous):

named x the width of the table the length would be x+7 the area of a rectangle is: A=w*l so A=x*(x+7)=78 x²+7x-78=0 delta=b²-4ac=7²-4*1*-78=361=19² x=(-7+19)/2 or (-7-19)/2 <= no possible x=12/2=6 width=6 length=13

OpenStudy (anonymous):

x(x + 7) = 78 x^2 + 7x -78 = 0 now factorise... (x-6)(x+13)= 0 x = 6 or -13 x cannot be equalto -13 as it is a length so x = 6 hence dimensions of table are 6 and (7+6) which is clearly 13